OpenStudy (anonymous):

solve each equation x-3/ x-2 - 1/x = x-3/x

jhonyy9 (jhonyy9):

x(x-3) -(x-2) =(x-2)(x-3) x^2 -3x -x +2 = x^2 -5x +6 -4x +2 = -5x +6 x=4